OpenStudy (anonymous):

how to take square root of 22801 number??

OpenStudy (anonymous):

just find a no. near to that of which you know sq root and then use differentiation..

OpenStudy (atlas):



OpenStudy (atlas):

STEP1.........Pair your numbers from the right

OpenStudy (atlas):

STEP2......Take the first number from the left (here it is 2) and think of a number whose square is less than or equal to the number Here it is 1

OpenStudy (atlas):

1*1 =1........I will write 1 on the left and 1 on the top .......>>>Also subtract 1 frm 2 (1 is the remainder)

OpenStudy (atlas):

Simultaneously add 1+1 =2

OpenStudy (atlas):

STEP 3.....Now bring down the next pair.........u get 128......>>>>>Think of a number such that 2? *? <=128...........here '?' is a single digit number

OpenStudy (atlas):

I found that 25*5 =125<128

OpenStudy (anonymous):

Thanks a lot @ atlas

OpenStudy (atlas):

Now again I write 5 on the top and add 25+5=30

OpenStudy (atlas):

continue the process till u get 0

hartnn (hartnn):

nice! a big round of applause for @atlas *claps*

OpenStudy (atlas):

haha.........thanks :)

OpenStudy (anonymous):

I really didnt get that method

OpenStudy (anonymous):

Can you please explain it

OpenStudy (atlas):

yeah sure

OpenStudy (anonymous):

similarly find the square root of 469225

OpenStudy (atlas):

you got it till what point

ganeshie8 (ganeshie8):

till pairing i got

OpenStudy (atlas):

6 -------------- 6 | 46 92 25 6*6 =36

OpenStudy (atlas):

6 -------------- 6 | 46 92 25 6*6 =36 +6 -36 I added 6+6 =12 on the left and subtracted 36 from 46 ... ------- 12 10

OpenStudy (atlas):

6? -------------- 6 | 46 92 25 6*6 =36 +6 -36 I added 6+6 =12 on the left and subtracted 36 from 46 ... ------- 12? 1092 I brought down next pair

OpenStudy (atlas):

Now think of a number ? which makes 12? X ? close to 1092

ganeshie8 (ganeshie8):

8

OpenStudy (anonymous):

8

OpenStudy (atlas):

right

OpenStudy (atlas):

68 -------------- 6 | 46 92 25 6*6 =36 +6 -36 I added 6+6 =12 on the left and subtracted 36 from 46 ... ------- 128 1092 I brought down next pair + 8 - 1024 ----------------------- 136 68

OpenStudy (atlas):

68? -------------- 6 | 46 92 25 6*6 =36 +6 -36 I added 6+6 =12 on the left and subtracted 36 from 46 ... ------- 128 1092 I brought down next pair + 8 - 1024 ----------------------- 136? 6825 i bring the next pair down and do the proces again

OpenStudy (anonymous):

1365*5=6825

OpenStudy (atlas):

till I get 0 in the remainder

OpenStudy (atlas):

right

OpenStudy (atlas):

you got it...........so the answer is 685

OpenStudy (anonymous):

great.... I got the process. seen it for the first time.

OpenStudy (anonymous):

what will we get if the number is not a perfect square??? the loop will never stop???

hartnn (hartnn):

food for thought 68*69 = 4692

ganeshie8 (ganeshie8):

looks like we're finding sqrt for 46 00 00 first next, subtracting 36 00 00 next, finding sqrt of 10 92 00 ... after that i lost a bit...

OpenStudy (atlas):

@ganeshie8 what we are doing is we are finding the maximum significant digit that can be in the square root of the given number.............then we find the next and so on

hartnn (hartnn):

there is this cool thing about perfect squares ending in 25, the last digit of square root is 5, ofcourse. the remaining digits, leaving 25 can be expressed in the form n(n+1) and the square root will be (n5) example, 4225---->6*7=42---->(65)^2 15625---->12*13=156---->(125)^2 incase you guys didn't know.... :)

OpenStudy (debbieg):

I've never seen this method either (or if I have, it was a very long time ago and I've forgotten it... lol). Fascinating. Thanks for the lessons @atlas . I want to study this a bit later when I have more time and really wrap my head around it. I get the mechanics, but I'd like to understand the theory around it better. :)

ganeshie8 (ganeshie8):

625 = 2(2+1), 5 = 25 cool trick ! @hartnn this works oly for perfect squares i think

hartnn (hartnn):

yes, only for perfect squares ending in 25

ganeshie8 (ganeshie8):

for 725, we cant write it as n(n+1)

hartnn (hartnn):

that trick is taken from a VEDIC math book. many such trick are there, which makes calculations much much easier...

OpenStudy (anonymous):

(n)*(n+1)*100+25 =(10n)^2 + 2* 10n * 5 +5^2 =(10n+5)^2

ganeshie8 (ganeshie8):

yup! i like VEDIC math :) i go thru once in a while.... somebody should write a book why/how VEDIC tricks works also (reasoning behind them)

hartnn (hartnn):

=(10n+5)^2 =(n5)^2

ganeshie8 (ganeshie8):

^^ saura u should write a companion book for VEDIC math... it sells im sure :)

hartnn (hartnn):

lol

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

@ganeshie8 did you get the theory behind what atlas did?

ganeshie8 (ganeshie8):

not yet, thats not obvious to me at all.. i gave up after trying for sometime. will get back to that in my evening ...

ganeshie8 (ganeshie8):

wbu saura ? it clicked to u?

OpenStudy (anonymous):

no.... I am thinking on it

OpenStudy (atlas):

I think it is pretty straight forward......the method does the algebraic manipulation at the backend: for the above example (100x + 10y +z)^2 = 469225 10000x^2 + 100y^2 +z^2 + 2000xy + 20yz + 200xz = 469225

OpenStudy (atlas):

we can find the value of MSB (x) which should be 6 here because 7 gives us a no. greater than RHS: So it becomes 100y^2 +z^2 +2000xy +20yz +200xz = 469225 - 360000

OpenStudy (atlas):

Again we add x+x (6+6) and keep it in ten's place.......i.e like 20x(120) and we think of a number y (8) that makes the number............20x+y(128) and we multiply this with y again making it 20xy + y^2

OpenStudy (atlas):

We are inherently multiplying this number 20x + y^2 with 100 because you see we are subtracting it with 1092(00).......so what we are doing is we are subtracting 2000xy + y^2 from RHS after getting the value of y

OpenStudy (atlas):

THis process continues